OpenStudy (anonymous):

1/4 c is less than or greater too 4 ?? help plz;-;

OpenStudy (anonymous):

1/4 is how much of one?

OpenStudy (anonymous):

Sorry gtg

OpenStudy (johnweldon1993):

\[\large \frac{1}{4}c \le 4\] So what number (when divided by 4) is less than or equal to 4? We can find that out by multiplying both sides of this equation by 4... \[\large c \le 16\]